Ozone, O₃, is produced by the reaction represented by the following equation: NO₂ + O₂ ⟶ NO + O₃ . What is the maximum amount of O₃ that can be produced when 56g NO₂ and 27g O₂ are reacted?

Your answer



1
Expert's answer
2021-05-28T00:56:02-0400

Molar Mass of NO2 = 46.0055

Molar Mass of O2 = 31.998

56/46.0055 = 1.217 moles

27/31.998 = 0.844 moles

= 1.217 + 0.844 = 2.061 moles

= 2.061 × 15.999 × 3

= 98.912g of O3